OpenStudy (anonymous):

How do i solve this? (1/5)^-1 + (-1/2)^-2

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ya i kno but how do i get 9

OpenStudy (anonymous):

(1/5)^-1 = (5/1)^1 = 5 (-1/2)^-2 = (-2/1)^2 = 4

OpenStudy (anonymous):

5+4 = 9

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

lol i guess but i m confused if its (-2/)^2 how come we get positive 4

OpenStudy (anonymous):

(-2) * (-2) = 4

OpenStudy (anonymous):

ohh ok and the powers wat happend to them?

OpenStudy (anonymous):

(-2)^2 = (-2) * (-2) That is the power of two. Meaning two of (-2)

OpenStudy (anonymous):

ahh ok aright thx a lot